# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v1

@@ Abstract
Distributed model predictive control (DMPC) has attracted extensive attention as it can explicitly handle system constraints and achieve optimal control in a decentralized manner. However, the deployment of DMPC strategies generally requires the sharing of sensitive data among subsystems, which may violate the privacy of participating systems. In this paper, we propose a differentially-private DMPC algorithm for linear discrete-time systems subject to coupled global constraints. Specifically, we first show that a conventional distributed dual gradient algorithm can be used to address the considered DMPC problem but cannot provide strong privacy preservation. Then, to protect privacy against the eavesdropper, we incorporate a differential-privacy noise injection mechanism into the DMPC framework and prove that the resulting distributed optimization algorithm can ensure both provable convergence to a global optimal solution and rigorous $\epsilon$-differential privacy. In addition, an implementation strategy of the DMPC is designed such that the recursive feasibility and stability of the closed-loop system are guaranteed. Numerical simulation results confirm the effectiveness of the proposed approach.
keywords:
Distributed Model Predictive Control , Privacy Preservation , Differential Privacy
journal: Automatica
1 Introduction
@@
Over the past decades, model predictive control (MPC) has enjoyed great success due to its ability to explicitly handle system constraints and guarantee prescribed control performance (Mayne, 2014; Christofides et al., 2013). MPC can be implemented either in a centralized or distributed manner. Centralized MPC relies on a central unit to process all system information and solve the online optimization problem, which often results in poor scalability and requires substantial computation power, especially for complex or large-scale systems. Consequently, distributed MPC (DMPC) has garnered much attention in recent years, offering many advantages of distributed systems and proving to be an effective tool for various applications, including vehicle platoons (Zheng et al., 2016), microgrids (Hans et al., 2018), and multi-robot systems (Luis et al., 2020).
@@
Based on the type of couplings between subsystems, DMPC studies can be roughly divided into three categories, i.e., coupling in cost functions, coupling in system dynamics, and coupling in constraints. In this paper, we focus on systems with coupled global constraints, which have many real-world applications (Bai et al., 2023; Michos et al., 2022). Several approaches have been proposed to guarantee the strict satisfaction of coupled constraints in a distributed manner. In Richards and How (2007), a sequential DMPC method is developed, which first divides the global problem into several local subproblems and then solves them sequentially in a given order. The satisfaction of coupled constraints is guaranteed through plan exchanges among subsystems. Trodden and Richards (2010) proposes a DMPC approach for flexible communication that can remove optimization order restrictions. Building on Richards and How (2007) and Trodden and Richards (2010), Trodden (2014) presents a parallel DMPC approach allowing simultaneous optimizations and maintaining flexible communication and parallel computation advantages. However, global optimality remains unclear in these approaches. In Wang and Ong (2017), a dual decomposition-based DMPC approach is designed to explicitly pursue global optimality. This approach transforms the dual problem of DMPC into a consensus optimization problem and then solves it using the distributed alternating direction multiplier method (ADMM). To improve the convergence speed when solving DMPC, a Nesterov-accelerated-gradient algorithm is utilized in Wang and Ong (2018). In addition, a push-sum dual gradient algorithm and a noisy ADMM algorithm are developed in Jin et al. (2020) and Li et al. (2021) to address the DMPC problem under time-varying directed communication network and communication noise, respectively.
@@
In the aforementioned methods (Wang and Ong, 2017, 2018; Jin et al., 2020; Li et al., 2021), distributed optimization algorithms are leveraged to address the DMPC problem, requiring each subsystem/agent to share explicit local information with neighboring subsystems/agents in order to satisfy coupled global constraints. Note that the shared messages often contain sensitive information, which raises significant concerns about privacy leakage. For instance, an eavesdropper could wiretap the communication channel and deduce privacy-sensitive information from exchanged messages. When DMPC is adopted in specific domains like smart grid or intelligent transportation, the disclosure of privacy-sensitive information may further pose safety risks and lead to economic losses. Considering the growing awareness of privacy and security, it is imperative to ensure privacy protection in DMPC. So far, few results are available on the privacy preservation of DMPC, but privacy-preserving approaches for distributed optimization have been well studied. For the latter, one typical technique is (partially) homomorphic encryption, which has been employed in Lu and Zhu (2018); Zhang and Wang (2018). The encryption-based methods use cryptography to mask privacy-sensitive information and can be directly extended for the privacy protection of DMPC (Zhao et al., 2023). However, this technique generally requires tedious encryption and decryption procedures, leading to huge overheads in both communication and computation. Another technique relies on spatially or temporally correlated noises/uncertainties (Gade and Vaidya, 2018; Lou et al., 2017; Gao et al., 2023), aiming to obscure information shared in distributed optimization. Due to the correlated nature of these noises/uncertainties, these approaches are typically vulnerable to adversaries with access to all messages shared in the network.
@@
With implementation simplicity and rigorous mathematical foundations, differential privacy (DP) has witnessed growing popularity, emerging as a de facto standard for privacy protection. In recent years, DP-based privacy methods have been introduced in distributed optimization by injecting DP noise to objective functions (Nozari et al., 2016) or exchanged information (Huang et al., 2015; Xiong et al., 2020; Ding et al., 2021). Nevertheless, the direct integration of persistent DP noise into existing algorithms inevitably compromises optimization accuracy, resulting in a fundamental trade-off between privacy and accuracy. It is crucial to note that extending DP-based privacy approaches to DMPC is not straightforward, as the compromise on optimization accuracy can deteriorate control performance and potentially lead to constraint violations.
@@
In this paper, a differentially-private DMPC algorithm is designed for linear discrete-time systems with coupled global constraints. We first demonstrate the need for privacy preservation by showcasing that a conventional distributed dual gradient algorithm for DMPC is vulnerable to eavesdropping attacks. A DP noise injection mechanism is then introduced into the distributed dual gradient algorithm, which obscures the private information exchanged among subsystems to prevent adversaries from inferring sensitive information. By leveraging the results in Wang (2023); Wang and Nedić (2024), a weakening factor sequence and a step-size sequence are carefully designed to effectively mitigate the influence of DP noise. Rigorous analysis shows that the proposed algorithm can ensure almost sure convergence to a global optimal solution and maintain $\epsilon$-differential privacy with a finite cumulative privacy budget. Aligned with the privacy-preserving distributed algorithm, we provide an implementation strategy for DMPC, ensuring the recursive feasibility and stability of the closed-loop system. Simulations are performed to validate the performance of the proposed scheme.
@@
The rest of the paper is organized as follows. In Section 2, the preliminaries of DMPC and differential privacy are introduced. In Section 3, a novel differentially-private distributed dual gradient algorithm is designed, and convergence analysis is conducted. In Section 4, the implementation strategy of DMPC is presented. Finally, a numerical study is given in Section 5, and concluding remarks are summarized in Section 6.
@@
Notations: $\mathbb{R}^{n}$ stands for the $n$-dimensional Euclidean space. Given two integers $a$ and $b$ ($a<b$), $\mathbb{Z}_{a}^{b}$ represents the set $\{a,a+1,\cdots,b\}$. $I_{n}$ denotes the identify matrix of dimension $n$. ${\bf 1}_{n}$ and ${\bf 0}_{n}$ represent the $n$-dimensional column vector with all entries equal to 1 and 0, respectively. We use $Q>(\geq)0$ to denote that $Q$ is a positive definite (semi-definite) matrix. $\|x\|$ and $\|x\|_{1}$ represent the standard Euclidean norm and the $L_{1}$ norm of a vector $x$, respectively. Moreover, $\|x\|_{Q}^{2}:=x^{\top}Qx$.
2 Problem Formulation and Preliminaries
2.1 Problem Description
@@
where $\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathcal{U}_{i}\subset\mathbb{R}^{m_{i}}$ are state and control input constraint sets, respectively. Moreover, all the subsystems are subject to $p$ global constraints described by
@@
Each linear discrete-time subsystem, i.e., $(A_{i},B_{i})$, is controllable. Additionally, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded and closed polytopes containing the origins as their inner point.
In this paper, we consider the same DMPC optimization problem as presented in Wang and Ong (2017, 2018); Jin et al. (2020); Li et al. (2021); Zhao et al. (2023). Specifically, based on (1)-(3), the DMPC optimization problem is formulated as
@@
where $N\in\mathbb{Z}_{>0}$ is the length of prediction horizon, $\tilde{x}_{i}(\ell|t)$ and $\tilde{u}_{i}(\ell|t)$ are the $\ell$th step-predicted state and control input at time instant $t$, respectively, $\tilde{{\bm{u}}}_{i}:=\{\tilde{u}_{i}(0|t),\cdots,\tilde{u}_{i}(N-1|t)\}$ stands for the predicted input sequence over the prediction horizon, and $Q_{i}>0$, $R_{i}>0$, and $P_{i}>0$ are weight matrices. For each subsystem $i$, $P_{i}$ is the solution of the algebraic Riccati equation
@@
where $K_{i}:=-(R_{i}+B_{i}^{\top}P_{i}B_{i})^{-1}B_{i}^{\top}P_{i}A_{i}$. The local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ in (4b) is defined as
$\tilde{\mathcal{U}_{i}}(x_{i}(t)):=\{\tilde{{\bm{u}}}_{i}\in\mathbb{R}^{m_{i}N}:\tilde{x}_{i}(\ell+1|t)=A_{i}\tilde{x}_{i}(\ell|t)+B_{i}\tilde{u}_{i}(\ell|t),$ (7)
$\tilde{x}_{i}(0|t)=x_{i}(t),\tilde{x}_{i}(\ell|t)\in\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\in\mathcal{U}_{i},\tilde{x}_{i}(N|t)\in\mathcal{X}_{i}^{f},\ell\in\mathbb{Z}_{0}^{N-1}\},$
@@
For more details about the tightening of constraint (3) and the construction of the terminal constraint set $\mathcal{X}_{i}^{f}$, please refer to Wang and Ong (2017).
@@
The communication network of $M$ subsystems is described by an interaction weight matrix $L=\{L_{ij}\}\in\mathbb{R}^{M\times M}$. Specifically, for each subsystem $i$, its neighbor set $\mathcal{N}_{i}$ is defined as the collections of subsystems $j$ such that subsystem $i$ and subsystem $j$ can directly communicate with each other. If $j\in\mathcal{N}_{i}$, then $L_{ij}>0$; otherwise, $L_{ij}=0$. We define $L_{ii}:=-\sum_{j\in\mathcal{N}_{i}}L_{ij}$ for all $i\in\mathbb{Z}_{1}^{M}$. Moreover, $L$ satisfies the following assumption:
@@
The matrix $L=\{L_{ij}\}$ is symmetric and satisfies ${\bf 1}_{M}^{\top}L={\bf 0}_{M}^{\top}$, $L{\bf 1}_{M}={\bf 0}_{M}$, and $\|I_{M}+L-\frac{{\bf 1}_{M}{\bf 1}_{M}^{\top}}{M}\|<1$.
Assumption 3 ensures that the communication graph induced by $L$ is connected, i.e., there is a path from any subsystem to any other subsystem.
2.2 Distributed Dual-Gradient Method
The Lagrangian function associated with the optimization problem in (4) is given by
@@ 2.2 Distributed Dual-Gradient Method
Based on Assumptions 1, 2 and the definition of DMPC problem, it can be concluded that the strong duality holds for (4), and the optimization problem (4) can be addressed by solving its dual problem (11). In addition, the Saddle-Point Theorem holds, i.e., given an optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$, we have
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda)\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda^{*}),\;\lambda\in\mathbb{R}^{Np}_{+},\;\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t)).$ (12)
A conventional approach to solving problem (11) is the distributed dual-gradient method (Falsone et al., 2017; Notarstefano et al., 2019). The core idea is to regard the Lagrangian multiplier (dual variable) $\lambda$ as a consensus variable and then $M$ subsystems address the optimization problem in a collective manner. Specifically, let each subsystem have a local copy $\lambda_{i}^{k}$ of the dual variable. $\Pi_{\mathbb{R}^{Np}_{+}}\left[\cdot\right]$ denotes Euclidean projection of a vector on the set +$\mathbb{R}^{Np}_{+}$, and $\gamma^{k}>0$ denotes the step-size. Then, the distributed dual-gradient method is summarized in Algorithm 1, and the overall DMPC implementation is detailed in Algorithm 2.

### v2

@@ Abstract
Distributed model predictive control (DMPC) has attracted extensive attention as it can explicitly handle system constraints and achieve optimal control in a decentralized manner. However, the deployment of DMPC strategies generally requires the sharing of sensitive data among subsystems, which may violate the privacy of participating systems. In this paper, we propose a differentially-private DMPC algorithm for linear discrete-time systems subject to coupled global constraints. Specifically, we first show that a conventional distributed dual gradient algorithm can be used to address the considered DMPC problem but cannot provide strong privacy preservation. Then, to protect privacy against the eavesdropper, we incorporate a differential-privacy noise injection mechanism into the DMPC framework and prove that the resulting distributed optimization algorithm can ensure both provable convergence to a global optimal solution and rigorous $\epsilon$-differential privacy. In addition, an implementation strategy of the DMPC is designed such that the recursive feasibility and stability of the closed-loop system are guaranteed. Simulation results are provided to demonstrate the effectiveness of the developed approach.
Index Terms:
Distributed model predictive control, privacy preservation, differential privacy.
I Introduction
@@
Over the past decades, model predictive control (MPC) has enjoyed great success due to its ability to explicitly handle system constraints and guarantee prescribed control performance [1, 2]. MPC can be implemented either in a centralized or distributed manner. Centralized MPC relies on a central unit to process all system information and solve the online optimization problem, which often results in poor scalability and requires substantial computation power, especially for complex or large-scale systems. Consequently, distributed MPC (DMPC) has garnered much attention in recent years, offering many advantages of distributed systems and proving to be an effective tool for various applications, including vehicle platoons [3], microgrids [4], and multi-robot systems [5].
@@
Based on the type of couplings between subsystems, DMPC studies can be roughly divided into three categories, i.e., coupling in cost functions, coupling in system dynamics, and coupling in constraints. In this paper, we focus on systems with coupled global constraints, which have many real-world applications [6, 7]. Several approaches have been proposed to guarantee the strict satisfaction of coupled constraints in a distributed manner. In [8], a sequential DMPC method is developed, which first divides the global problem into several local subproblems and then solves them sequentially in a given order. The satisfaction of coupled constraints is guaranteed through plan exchanges among subsystems. [9] proposes a DMPC approach for flexible communication that can remove optimization order restrictions. Building on [8] and [9], [10] presents a parallel DMPC approach allowing simultaneous optimizations and maintaining flexible communication and parallel computation advantages. However, global optimality remains unclear in these approaches. In [11], a dual decomposition-based DMPC approach is designed to explicitly pursue global optimality. This approach transforms the dual problem of DMPC into a consensus optimization problem and then solves it using the distributed alternating direction multiplier method (ADMM). To improve the convergence speed when solving DMPC, a Nesterov-accelerated-gradient algorithm is utilized in [12]. In addition, a push-sum dual gradient algorithm and a noisy ADMM algorithm are developed in [13] and [14] to address the DMPC problem under time-varying directed communication network and communication noise, respectively.
@@
In the aforementioned methods [11, 12, 13, 14], distributed optimization algorithms are leveraged to address the DMPC problem, requiring each subsystem/agent to share explicit local information with neighboring subsystems/agents in order to satisfy coupled global constraints. Note that the shared messages often contain sensitive information, which raises significant concerns about privacy leakage. For instance, an eavesdropper could wiretap the communication channel and deduce privacy-sensitive information from exchanged messages. When DMPC is adopted in specific domains like smart grid or intelligent transportation, the disclosure of privacy-sensitive information may further pose safety risks and lead to economic losses. Considering the growing awareness of privacy and security, it is imperative to ensure privacy protection in DMPC. So far, few results are available on the privacy preservation of DMPC, but privacy-preserving approaches for distributed optimization have been well studied. For the latter, one typical technique is (partially) homomorphic encryption, which has been utilized in existing works such as [15, 16]. The encryption-based approaches use cryptography to conceal privacy-sensitive information and can be directly extended for the privacy protection of DMPC [17]. However, this technique generally requires tedious encryption and decryption procedures, leading to huge overheads in both communication and computation. Another technique relies on spatially or temporally correlated noises/uncertainties [18, 19, 20], aiming to obscure information shared in distributed optimization. Due to the correlated nature of these noises/uncertainties, these approaches are typically vulnerable to adversaries with access to all messages shared in the communication network.
@@
With implementation simplicity and rigorous mathematical foundations, differential privacy (DP) has witnessed growing popularity, emerging as a de facto standard for privacy protection. In recent years, DP-based privacy methods have been introduced in distributed optimization by integrating DP noise into objective functions [21] or exchanged information [22, 23, 24]. Nevertheless, the direct injection of persistent DP noise to existing algorithms inevitably compromises optimization performance, resulting in an inherent trade-off between accuracy and privacy. It is crucial to note that extending DP-based privacy approaches to DMPC is not straightforward, as the compromise on optimization accuracy can deteriorate control performance and potentially lead to constraint violations.
@@
In this paper, a differentially-private DMPC algorithm is designed for linear discrete-time systems with coupled global constraints. We first demonstrate the need for privacy preservation by showcasing that a conventional distributed dual gradient algorithm for DMPC is vulnerable to eavesdropping attacks. A DP noise injection mechanism is then introduced into the distributed dual gradient algorithm, which obscures the private information exchanged among subsystems to prevent adversaries from inferring sensitive information. By leveraging the results in [25, 26], a weakening factor sequence and a step-size sequence are carefully designed to effectively mitigate the influence of DP noise. Rigorous analysis shows that the proposed algorithm can ensure almost sure convergence to a global optimal solution and maintain $\epsilon$-differential privacy with a finite cumulative privacy budget. Aligned with the privacy-preserving distributed algorithm, we provide an implementation strategy for DMPC, ensuring the recursive feasibility and stability of the closed-loop system. Simulations are performed to validate the performance of the proposed scheme.
@@
The rest of this paper is organized as follows. In Section II, the preliminaries of DMPC and differential privacy are introduced. In Section III, a new differentially-private distributed dual gradient algorithm is developed, and convergence analysis is conducted. Section IV presents the implementation strategy of DMPC. Finally, a numerical study is given in Section V, and concluding remarks are summarized in Section VI.
@@
Notations: $\mathbb{R}^{n}$ stands for the $n$-dimensional Euclidean space. Given two integers $a$ and $b$ ($a<b$), $\mathbb{Z}_{a}^{b}$ represents the set $\{a,a+1,\cdots,b\}$. $I_{n}$ denotes the identify matrix of dimension $n$. ${\bf 1}_{n}$ and ${\bf 0}_{n}$ represent the $n$-dimensional column vector with all entries being 1 and 0, respectively. We use $Q>(\geq)0$ to denote that $Q$ is a positive definite (semi-definite) matrix. $\|x\|$ and $\|x\|_{1}$ represent the standard Euclidean norm and the $L_{1}$ norm of a vector $x$, respectively. Moreover, $\|x\|_{Q}^{2}:=x^{\top}Qx$.
II Problem Formulation and Preliminaries
II-A Problem Description
@@
where $\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathcal{U}_{i}\subset\mathbb{R}^{m_{i}}$ denote state and control input constraint sets, respectively. Moreover, all the subsystems are subject to $p$ global constraints described by
@@
Each linear discrete-time subsystem, i.e., $(A_{i},B_{i})$, is controllable. Additionally, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded and closed polytopes which contain the origins as their inner point.
In this paper, we consider the same DMPC problem as presented in [11, 12, 13, 14, 17]. Specifically, based on (1)-(3), the DMPC problem is formulated as
@@
where $N\in\mathbb{Z}_{>0}$ is the length of prediction horizon, $\tilde{x}_{i}(\ell|t)$ and $\tilde{u}_{i}(\ell|t)$ are the $\ell$th step predicted state and control input at time instant $t$, respectively, $\tilde{{\bm{u}}}_{i}:=\{\tilde{u}_{i}(0|t),\cdots,\tilde{u}_{i}(N-1|t)\}$ stands for the predicted input sequence over the prediction horizon, and $Q_{i}>0$, $R_{i}>0$, and $P_{i}>0$ are weight matrices. For each subsystem $i$, $P_{i}$ is the solution of the following algebraic Riccati equation:
@@
where $K_{i}:=-(R_{i}+B_{i}^{\top}P_{i}B_{i})^{-1}B_{i}^{\top}P_{i}A_{i}$. The local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ in (4b) is formulated as
$\tilde{\mathcal{U}_{i}}(x_{i}(t)):=\{\tilde{{\bm{u}}}_{i}\in\mathbb{R}^{m_{i}N}:$ (7)
$\tilde{x}_{i}(\ell+1|t)=A_{i}\tilde{x}_{i}(\ell|t)+B_{i}\tilde{u}_{i}(\ell|t),\tilde{x}_{i}(0|t)=x_{i}(t),$
$\tilde{x}_{i}(\ell|t)\in\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\in\mathcal{U}_{i},\tilde{x}_{i}(N|t)\in\mathcal{X}_{i}^{f},\ell\in\mathbb{Z}_{0}^{N-1}\},$
@@
For more details about the tightening of constraint (3) and the construction of the terminal constraint set $\mathcal{X}_{i}^{f}$, please refer to [11].
@@
The communication network of $M$ subsystems is described by an interaction weight matrix $L=\{L_{ij}\}\in\mathbb{R}^{M\times M}$. Specifically, for each subsystem $i$, the neighbor set $\mathcal{N}_{i}$ consists of all subsystems $j$ that can directly communicate with subsystem $i$. If $j\in\mathcal{N}_{i}$, then $L_{ij}>0$; otherwise, $L_{ij}=0$. We define $L_{ii}:=-\sum_{j\in\mathcal{N}_{i}}L_{ij}$ for all $i\in\mathbb{Z}_{1}^{M}$. Moreover, $L$ satisfies the following assumption:
@@
The interaction weight matrix $L=\{L_{ij}\}$ is symmetric and satisfies ${\bf 1}_{M}^{\top}L={\bf 0}_{M}^{\top}$, $L{\bf 1}_{M}={\bf 0}_{M}$, and $\|I_{M}+L-\frac{{\bf 1}_{M}{\bf 1}_{M}^{\top}}{M}\|<1$.
Assumption 3 guarantees that the communication network described by $L$ is connected, meaning that there exists a path from any subsystem to any other subsystem.
II-B Distributed Dual-Gradient Method
The Lagrangian function corresponding to the optimization problem in (4) is given by
@@ II-B Distributed Dual-Gradient Method
Based on Assumptions 1, 2 and the definition of DMPC problem, it can be concluded that the strong duality holds for (4), and the optimization problem (4) can be addressed by solving its dual problem (11). In addition, the Saddle-Point Theorem holds, i.e., given an optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$, the following relationship holds for any $\lambda\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$:
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda)\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda^{*}).$ (12)
A conventional approach to solving problem (11) is the distributed dual-gradient method [27, 28]. The core idea is to regard the Lagrangian multiplier (dual variable) $\lambda$ as a consensus variable and then $M$ subsystems address the optimization problem in a collective manner. Specifically, let each subsystem have a local copy $\lambda_{i}^{k}$ of the dual variable. $\Pi_{\mathbb{R}^{Np}_{+}}\left[\cdot\right]$ denotes Euclidean projection of a vector on the set +$\mathbb{R}^{Np}_{+}$, and $\gamma^{k}>0$ denotes the step-size. Then, the distributed dual-gradient method is summarized in Algorithm 1, and the overall DMPC implementation is detailed in Algorithm 2.